At a price of $5, quantity demanded is 25 units, while at a price of $4, quantity demanded is 35 units. What is the price elasticity of demand when price falls from $5 to $4?

a

0.5

b

2.5

c

1.5

d

0.2
Answer #1

Price elasticity of demand in the market = % change in the demand / % change in the price.

% change in demand = 25-35 / 25 + 35

= -0.17

% change in price = 5 -4 / 5+4

= 1 / 9 = 0.11

Elasticity = 0.17 / 0.11 = 1.5. The answer is "C".